Prove that the pressure exerted by the gas is directy proportional to average transilated kinetic energy of the gas molecules

Grade:12th pass

1 Answers

Susmita
5 years ago
Ideal gas equation
PV=mnc2/3=(mnc2/2)×(2/3)=Em×(2/3)
where mean kinetic energy Em=(1/2)mnc2..................................................................................